# Universal Hitchin moduli spaces

### v1

@@ 4.2. Universal moduli space of the coupled harmonic equations
Assume that $\operatorname{ker}\mathcal{L}=\{0\}$. Then, there exists an orthogonal decomposition (5.20) for the pairing $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$. Consequently, for any element $v\in\mathcal{S}^{1}$ there exists a unique $\Pi v\in\operatorname{Im}\;\widetilde{\mathbf{P}}$ such that $(\dot{J},a,\dot{\psi})=v-\Pi v$ solves the linear equations
$\begin{split}\alpha d_{A}(Ja)-\alpha[\dot{\psi},J\psi]-\alpha d_{A}(\psi(J\dot{J}))&=0\\ \varepsilon\delta S(J\dot{J})+\alpha*d\left(B(\Lambda_{\omega}d_{A}\dot{\psi},*\psi)-B(\Lambda_{\omega}F_{A},*a)+B(\Lambda_{\omega}F_{A},\psi(J\dot{J}))\right)&=0.\end{split}$
@@
If $v\in\operatorname{Im}\;\widetilde{\mathbf{P}}\cap(\operatorname{Im}\;\widetilde{\mathbf{P}})^{\perp_{\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}}}$, then $v=\widetilde{\mathbf{P}}(y)$ for $y\in\mathcal{S}^{0}$. But then $\widetilde{\mathbf{P}}^{*}\circ\widetilde{\mathbf{P}}(y)=0$ and, by $\operatorname{ker}\mathcal{L}=\{0\}$, $v=0$. Thus
@@
But by Proposition 4.7 and condition $\operatorname{ker}\mathcal{L}=\{0\}$, $\widetilde{\mathbf{P}}^{*}\circ{\bf\widetilde{P}}$ is surjective. Then, by elliptic regularity, one can solve (4.15) for $y\in\mathcal{S}^{0}$. The orthogonal decomposition follows. The last statement of the lemma comes from the expression of $\bf{\widetilde{P}}^{*}$ in Lemma 5.14. ∎
@@
Assume $\operatorname{ker}\mathcal{L}=\{0\}$ and $\alpha\neq 0$. Then, the inclusion $\mathcal{H}^{1}(\mathcal{S}^{*})\subset\ker\bf{L}^{\varepsilon}_{\alpha}$ induces an isomorphism
$\mathcal{H}^{1}(\mathcal{S}^{*})\simeq H^{1}(\mathcal{S}^{*}).$
@@
(4.16) $\begin{split}d_{A}a-[\dot{\psi},\psi]&=0,\\ d_{A}\dot{\psi}+[a,\psi]&=0,\\ d_{A}(J\dot{\psi})+[a,J\psi]-d_{A}(\psi(\dot{J}))&=0,\\ d_{A}(Ja)-[\dot{\psi},J\psi]-d_{A}(\psi(J\dot{J}))&=0\\ \varepsilon\delta S(\dot{J})-\alpha*d\left(B(\Lambda_{\omega}d_{A}a,*\psi)+B(\Lambda_{\omega}F_{A},*\dot{\psi})-B(\Lambda_{\omega}F_{A},\psi(\dot{J}))\right)&=0,\\ \varepsilon\delta S(J\dot{J})-\alpha*d\left(-B(\Lambda_{\omega}d_{A}\dot{\psi},*\psi)+B(\Lambda_{\omega}F_{A},*a)-B(\Lambda_{\omega}F_{A},\psi(J\dot{J}))\right)&=0,\end{split}$
@@
$\mathcal{U}^{*}_{\alpha,\varepsilon}=\{[(J,A,\psi)]\;|\;\operatorname{ker}\mathcal{L}=\{0\}\}\subset\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}.$
@@
$\Phi=\varepsilon\nu{J}+\frac{\alpha}{2}\|\psi\|^{2}_{L^{2}}$
and $\nu{J}$ is induced by the global Kähler potential in the space of complex structures compatible with the orientation $\curly{J}$ (see [F, Section 4]).
@@
The fact that $\mathcal{U}^{*}_{\alpha,\varepsilon}$ is open follows from upper semicontinuity of the dimension of the kernel for elliptic operators. Given now a smooth point $(J,A,\psi)]\in\mathcal{U}^{*}_{\alpha,\varepsilon}$, the tangent space is identified with $H^{1}(\mathcal{S}^{*})$ and inherits a complex structure by Proposition 4.9, using that $\mathbb{J}$ in (3.7) preserves (4.16). The existence the of the symmetric tensor $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ is a direct consequence of Lemma 3.4 and Proposition 3.9, while the listed signature properties follow from Lemma 3.8. The formula for the Kähler potential is a direct consequence of (3.13), while a explicit formula for $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ follows from Lemma 3.4. ∎
@@
(4.17) $\begin{split}\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v,v)&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2}))\\ &+\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge J(\psi(\dot{J}_{2}))),\end{split}$
@@ 4.3. Comparison with $\boldsymbol{\mathcal{U}^{\operatorname{Flat}}(G)}$ and existence
This aim of this section is twofold. Firstly, we establish a comparison between the moduli spaces $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ and the universal moduli space of flat $G$-connections $\mathcal{U}^{\operatorname{Flat}}(G)$ constructed in Section 4.1. As we will see, for any $\alpha>0$ and $\varepsilon\neq 0$, Theorem 4.10 induces a natural holomorphic map
$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\supset\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)$
and hence a holomorphic map into Teichmuller space $\mathcal{T}$. Secondly, we will prove that for genus of the surface $\Sigma$ bigger than zero, the open $\mathcal{U}^{*}_{\alpha,\varepsilon}\subset\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ is non-empty for sufficiently small values of $\alpha$.
@@
The tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{\operatorname{Flat}}(G)$ can be formally identified with the cohomology of the complex of linear differential operators
(4.19) ${(\mathcal{C}^{*})}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{C}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{P}^{c}}$${\mathcal{C}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\bf L^{c}}}$${\mathcal{C}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
@@
$\mathcal{C}^{0}=\operatorname{Lie}\widetilde{\curly{G}}\cong\Omega^{0}(T\Sigma)\oplus\Omega^{0}(X,E_{G}(\mathfrak{g})),\qquad\mathcal{C}^{1}=\mathcal{S}^{1},\qquad\mathcal{C}^{2}=\Omega^{2}(X,E_{G}(\mathfrak{g}))$
@@
$(\dot{J},a,\dot{\psi})=\sigma_{\mathbf{P}^{c}}(v)(y,u_{0}+i(u_{1}^{\prime}-\psi(y))).$
@@
(4.20) $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)\colon[(J,A,\psi)]\mapsto[(J,A+i\Psi)].$
@@
(4.21) ${H^{1}(\mathcal{S}^{*})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\mathcal{C}^{*}).}$
@@
$H^{1}(\mathcal{S}^{*})\simeq\mathcal{H}^{1}(\mathcal{S}^{*}),$
where the right hand side is given by solutions of the gauge-fixed system of linear equations (4.16). Then, the map
$\mathcal{H}^{1}(\mathcal{S}^{*})\to H^{1}(\mathcal{C}^{*})\colon[(\dot{J},a,\dot{\psi})]\mapsto[(\dot{J},a+i\dot{\psi})]$
@@
(4.22) $\begin{split}d_{A_{h}}^{*}\psi_{h}&=0\\ \varepsilon S_{g}-\alpha*d\left(B\left(\Lambda_{\omega}F_{A},*\psi_{h}\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V}\end{split}$
@@
$\mathbf{Q}_{\alpha}\colon C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(X,E_{K}(\mathfrak{k}))\to C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{2}(X,E_{K}(\mathfrak{k}))$
@@
$\mathbf{Q}_{\alpha}(f,u)=(S_{g_{f}}-\varepsilon^{-1}\alpha*d\left(B\left(\Lambda_{\omega_{f}}F_{A_{h_{u}}},*\psi_{h_{u}}\right)\right)-2\pi\chi(\Sigma)V^{-1},\operatorname{Ad}(e^{-iu})d_{A_{h_{u}}}J\psi_{h_{u}})$
@@
$\mathcal{N}^{\alpha}=(\varepsilon^{-1}\widetilde{\mathbf{P}}_{1}^{*},\alpha^{-1}\widetilde{\mathbf{P}}_{2}^{*})\circ\widetilde{\mathbf{P}}\colon\mathcal{S}^{0}\to\mathcal{S}^{0}.$
@@
$\mathcal{N}^{0}=\mathcal{N}^{0}_{0}\oplus\mathcal{N}^{0}_{1}$
@@
The last part of the statement follows from that fact that (5.25) is induced by a $C^{\infty}$-Frechet map, combined with Lemma 5.21.

### v2

@@ 4.2. Universal moduli space of the coupled harmonic equations
Assume that $\ker\mathcal{L}=\{0\}$. Then, there exists an orthogonal decomposition (5.20) for the pairing $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$. Consequently, for any element $v\in\mathcal{S}^{1}$ there exists a unique $\Pi v\in\operatorname{Im}\;\widetilde{\mathbf{P}}$ such that $(\dot{J},a,\dot{\psi})=v-\Pi v$ solves the linear equations
$\begin{split}\alpha d_{A}(Ja)-\alpha[\dot{\psi},J\psi]-\alpha d_{A}(\psi(J\dot{J}))&=0,\\ \varepsilon\delta S(J\dot{J})+\alpha*d\left(B(\Lambda_{\omega}d_{A}\dot{\psi},*\psi)-B(\Lambda_{\omega}F_{A},*a)+B(\Lambda_{\omega}F_{A},\psi(J\dot{J}))\right)&=0.\end{split}$
@@
If $v\in\operatorname{Im}\;\widetilde{\mathbf{P}}\cap(\operatorname{Im}\;\widetilde{\mathbf{P}})^{\perp_{\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}}}$, then $v=\widetilde{\mathbf{P}}(y)$ for $y\in\mathcal{S}^{0}$. But then $\widetilde{\mathbf{P}}^{*}\circ\widetilde{\mathbf{P}}(y)=0$ and, by $\ker\mathcal{L}=\{0\}$, $v=0$. Thus
@@
But by Proposition 4.7 and condition $\ker\mathcal{L}=\{0\}$, $\widetilde{\mathbf{P}}^{*}\circ{\bf\widetilde{P}}$ is surjective. Then, by elliptic regularity, one can solve (4.15) for $y\in\mathcal{S}^{0}$. The orthogonal decomposition follows. The last statement of the lemma comes from the expression of $\bf{\widetilde{P}}^{*}$ in Lemma 5.14. ∎
@@
Assume $\ker\mathcal{L}=\{0\}$ and $\alpha\neq 0$. Then, the inclusion $\mathcal{H}^{1}(\mathcal{S}^{*})\subset\ker\bf{L}^{\varepsilon}_{\alpha}$ induces an isomorphism
$\mathcal{H}^{1}(\mathcal{S}^{*})\cong H^{1}(\mathcal{S}^{*}).$
@@
(4.16) $\begin{split}d_{A}a-[\dot{\psi},\psi]&=0,\\ d_{A}\dot{\psi}+[a,\psi]&=0,\\ d_{A}(J\dot{\psi})+[a,J\psi]-d_{A}(\psi(\dot{J}))&=0,\\ d_{A}(Ja)-[\dot{\psi},J\psi]-d_{A}(\psi(J\dot{J}))&=0\\ \varepsilon\delta S(\dot{J})-\alpha*d\left(B(\Lambda_{\omega}d_{A}a,*\psi)+B(\Lambda_{\omega}F_{A},*\dot{\psi})-B(\Lambda_{\omega}F_{A},\psi(\dot{J}))\right)&=0,\\ \varepsilon\delta S(J\dot{J})-\alpha*d\left(-B(\Lambda_{\omega}d_{A}\dot{\psi},*\psi)+B(\Lambda_{\omega}F_{A},*a)-B(\Lambda_{\omega}F_{A},\psi(J\dot{J}))\right)&=0.\end{split}$
@@
$\mathcal{U}^{*}_{\alpha,\varepsilon}=\{[(J,A,\psi)]\;|\;\ker\mathcal{L}=\{0\}\}\subset\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}.$
@@
$\Phi=\varepsilon\nu{J}+\frac{\alpha}{2}\|\psi\|^{2}_{L^{2}},$
and $\nu{J}$ is induced by the global Kähler potential in the space of complex structures compatible with the orientation $\curly{J}$ (see [Fu, Section 4]).
@@
The fact that $\mathcal{U}^{*}_{\alpha,\varepsilon}$ is open follows from upper semicontinuity of the dimension of the kernel for elliptic operators. Given now a smooth point $(J,A,\psi)]\in\mathcal{U}^{*}_{\alpha,\varepsilon}$, the tangent space is identified with $H^{1}(\mathcal{S}^{*})$ and inherits a complex structure by Proposition 4.9, using that $\mathbb{J}$ in (3.7) preserves (4.16). The existence of the symmetric tensor $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ is a direct consequence of Lemma 3.4 and Proposition 3.9, while the listed signature properties follow from Lemma 3.8. The formula for the Kähler potential is a direct consequence of (3.9), while an explicit formula for $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ follows from Lemma 3.4. ∎
@@
(4.17) $\begin{split}\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &+\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge J(\psi(\dot{J}_{2}))).\end{split}$
@@ 4.3. Comparison with $\boldsymbol{\mathcal{U}^{\operatorname{Flat}}(G)}$ and existence
The aim of this section is twofold. Firstly, we establish a comparison between the moduli spaces $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ and the universal moduli space of flat $G$-connections $\mathcal{U}^{\operatorname{Flat}}(G)$ constructed in Section 4.1. As we will see, for any $\alpha>0$ and $\varepsilon\neq 0$, Theorem 4.10 induces a natural holomorphic map
$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\supset\mathcal{U}^{*}_{\alpha,\varepsilon}\longrightarrow\mathcal{U}^{\operatorname{Flat}}(G),$
and hence a holomorphic map into Teichmüller space $\mathcal{T}$. Secondly, we will prove that for genus of the surface $\Sigma$ bigger than zero, the open $\mathcal{U}^{*}_{\alpha,\varepsilon}\subset\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ is non-empty for sufficiently small values of $\alpha$.
@@
The tangent space $T_{\left[(J,A,\psi)\right]}\mathcal{U}^{\operatorname{Flat}}(G)$ can be formally identified with the cohomology of the complex of linear differential operators
[…]
@@
$\mathcal{C}^{0}=\operatorname{Lie}\widetilde{\curly{G}}\cong\Omega^{0}(T\Sigma)\oplus\Omega^{0}(\Sigma,E_{G}(\mathfrak{g})),\qquad\mathcal{C}^{1}=\mathcal{S}^{1},\qquad\mathcal{C}^{2}=\Omega^{2}(\Sigma,E_{G}(\mathfrak{g})),$
@@
$(\dot{J},a,\dot{\psi})=\sigma_{\mathbf{P}^{c}}(v)(y,u_{0}+i(u_{1}^{\prime}-\psi(y))).\qed$
@@
(4.20) $\begin{array}[]{rcl}\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}&\longrightarrow&\mathcal{U}^{\operatorname{Flat}}(G)\\ \left[(J,A,\psi)\right]&\longmapsto&\left[(J,A+i\psi)\right].\end{array}$
@@
(4.21) $H^{1}(\mathcal{S}^{*})$$H^{1}(\mathcal{C}^{*}),$}
@@
$H^{1}(\mathcal{S}^{*})\cong\mathcal{H}^{1}(\mathcal{S}^{*}),$
where the right-hand side is given by solutions of the gauge-fixed system of linear equations (4.16). Then, the map
$\begin{array}[]{ccl}\mathcal{H}^{1}(\mathcal{S}^{*})&\longrightarrow&H^{1}(\mathcal{C}^{*})\\ \left[(\dot{J},a,\dot{\psi})\right]&\longmapsto&\left[(\dot{J},a+i\dot{\psi})\right]\end{array}$
@@
(4.22) $\begin{split}d_{A_{h}}^{*}\psi_{h}&=0,\\ \varepsilon S_{g}-\alpha*d\left(B\left(\Lambda_{\omega}F_{A},*\psi_{h}\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V},\end{split}$
@@
$\mathbf{Q}_{\alpha}\colon C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(\Sigma,E_{K}(\mathfrak{k}))\longrightarrow C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{2}(\Sigma,E_{K}(\mathfrak{k}))$
@@
$\mathbf{Q}_{\alpha}(f,u)=(S_{g_{f}}-\varepsilon^{-1}\alpha*d\left(B\left(\Lambda_{\omega_{f}}F_{A_{h_{u}}},*\psi_{h_{u}}\right)\right)-2\pi\chi(\Sigma)V^{-1},\operatorname{Ad}(e^{-iu})d_{A_{h_{u}}}J\psi_{h_{u}}),$
@@
$\mathcal{N}^{\alpha}=(\varepsilon^{-1}\widetilde{\mathbf{P}}_{1}^{*},\alpha^{-1}\widetilde{\mathbf{P}}_{2}^{*})\circ\widetilde{\mathbf{P}}\colon\mathcal{S}^{0}\longrightarrow\mathcal{S}^{0}.$
@@
$\mathcal{N}^{0}=\mathcal{N}^{0}_{0}\oplus\mathcal{N}^{0}_{1},$
@@
The last part of the statement follows from that fact that (4.20) is induced by a $C^{\infty}$-Frechet map, combined with Lemma 5.21. ∎